OpenStudy (anonymous):

How do I prove this: (secx-cosx)(cscx-sinx)=tanx/(1+tan^2x)

OpenStudy (anonymous):

left side: (1-cos^2x)/cosx * (1-sin^2x)/sinx (sin^2xcos^2x)/(sinxcosx) sinxcosx is that right?

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

you have a calculater?

OpenStudy (anonymous):

with the sin cos

OpenStudy (anonymous):

yeah i do, but imnot suppose to use a calculator for this

OpenStudy (mertsj):

\[(1/\cos-\cos)(1/\sin-\sin)=[ (1-\cos^2)/\cos][(1-\sin^2)/\sin]=(\sin^2/\cos)(\cos^2/\sin)\]